### v4

# Hidden Minima in Two-Layer ReLU Networks

Abstract.
We consider the optimization problem associated with training two-layer ReLU networks with $d$ inputs under the squared loss, where the labels are generated by a target network. Recent work has identified two distinct classes of infinite families of minima: one whose training loss vanishes in the high-dimensional limit, and another whose loss remains bounded away from zero. The latter family is empirically avoided by stochastic gradient descent, hence hidden, motivating the search for analytic criteria that distinguish hidden from non-hidden minima. A key challenge is that prior analyses have shown the Hessian spectra at hidden and non-hidden minima to coincide up to terms of order $O(d^{-1/2})$, seemingly limiting the discriminative power of spectral methods. We therefore take a different route, studying instead certain curves along which the loss is locally minimized. Our main result shows that arcs emanating from hidden minima exhibit distinctive structural and symmetry properties, arising precisely from $\Omega(d^{-1/2})$ eigenvalue contributions that are absent from earlier analyses.
1. Introduction
An outstanding question in deep learning (DL) concerns the ability of simple gradient-based methods to successfully train neural networks despite the nonconvexity of the associated optimization problems. Indeed, nonconvex optimization landscapes may have spurious (i.e., non-global local) minima with large basins of attraction and this can cause a complete failure of these methods. Our understanding of the nature by which nonconvex problems associated with artificial neural networks differ from computationally hard ones is currently limited. In view of the complexity exhibited by contemporary networks and the absence of suitable analytic tools, much recent research has focused on two-layer ReLU networks as a realistic starting point for a theoretical study, e.g., [15, 18, 44, 35, 24, 47, 40]. The two-layer networks considered were typically of the form:
$f(\mathbf{x};W,\boldsymbol{a})\coloneqq\boldsymbol{a}^{\top}\!\sigma(W\mathbf{x}),\quad~W\in M(k,d),~\boldsymbol{a}\in\mathbb{R}^{k},$ (1.1)
where $\sigma(t)\coloneqq\max\{0,t\}$ is the ReLU function acting entrywise and $M(k,d)$ denotes the space of $k\times d$ matrices. To isolate the study of optimization-related obstructions on account of nonconvexity from those pertaining to the expressive power of two-layer networks, data has been often assumed to be fully realizable. This was further motivated by hardness results indicating a strict barrier inherent to the explanatory power of distribution-free approaches operating in complete generality [13, 16, 43]. For the squared loss, the resulting highly nonconvex expected loss is
$\mathcal{L}(W,\boldsymbol{\alpha})\coloneqq\frac{1}{2}\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}\left[\Big(f(\mathbf{x};W,\boldsymbol{a})-f(\mathbf{x};T,\boldsymbol{b})\Big)^{2}\right],$ (1.2)
where $\mathcal{D}$ denotes a probability distribution over the input space, $W\in M(k,d)$, $\boldsymbol{a}\in\mathbb{R}^{k}$ denote the optimization variables, and $T\in M(d,d)$, $\boldsymbol{b}\in\mathbb{R}^{d}$ are fixed parameters. Recently, it was found [5] that the symmetry of spurious minima of $\mathcal{L}$} break the symmetry of global minima under various choices of data distributions (formal terms are given later). Here, for concreteness, we focus on the $d$-variate normal distribution for inputs [20, 51, 32, 46, 23, 12].
Remark 1.
In [3], the theory of distributions is adopted as a central analytical framework for the rigorous treatment of the nonsmoothness inherent in ReLU networks. This approach applies to a broad class of data distributions given by Radon measures, that is, finite Borel measures of total mass one, and in particular includes the normal distribution as is the case in the present work. It furnishes a unified and mathematically precise foundation for differentiation with respect to network parameters, encompassing, in particular, bias terms (see the concluding remark in Section 5 below). Within this framework, techniques from geometric measure theory provide a refined structural characterization of the gradient, which is viewed as a vector-valued function of bounded variation, as well as of the Hessian, now represented in the distributional sense as a matrix-valued Radon measure.
Using ideas based on symmetry breaking, techniques from representation theory and real algebraic geometry were employed to construct infinite families of critical points represented by Puiseux series in $d^{-1}$ and so obtain sharp estimates for the loss and the Hessian spectrum holding for finite, arbitrarily large, dimensionality [8, 6, 7]. We refer to [4, 11] for similar analyses of tensor decomposition problems. The analytic results were used to investigate some of the key foundational phenomena occurring in DL. For example, the phenomenon of extremely skewed spectrum of the Hessian observed for large-scale trained networks [14, 31, 41, 42] was established, analytically, for families of minima and for arbitrarily large dimensionality. Other results concerned, for example, a long standing debate regarding whether some notion of local curvature can be used to explain generalization [27, 29, 28, 17, 19], ruling out, in the setting considered, notions of ‘flatness’ relying exclusively on the Hessian spectrum.
A particular phenomenon concerning two types of infinite families of critical points: type I and type II, was as follows. Despite type I and type II critical points being provably local minima for (essentially) all $d\in\mathbb{N}$, empirically, the former is never detected by standard gradient-based optimization methods initialized using, e.g., Xavier initialization. Thus, henceforth, type I minima shall also be referred to as hidden minima. This favoring of type II minima over type I reflects a bias of optimization methods towards minima of reduced loss. Indeed, while the loss at type II minima converges to zero as $d$ increases ($O(1/d)$ for all known type II families), at type I minima the loss remains bounded away from zero. In particular, any hidden minimum is necessarily spurious (but not vice verse. Type II spurious minima exist and are detected by gradient-based methods). Of course, global minima of $\mathcal{L}$}, loss zero, are type II.
Eigenvalue $\frac{\pi-2}{4\pi}$ $\frac{1}{4}$ $\frac{\pi+2}{4\pi}$ $\frac{d}{2\pi}+\frac{-\pi^{2}-4+6\pi}{4\pi\left(2-\pi\right)}$ $\frac{d}{4}+\frac{-\pi^{2}-2\pi+4}{4\pi\left(2-\pi\right)}$ $\frac{d}{4}+\frac{1}{4}$
Multiplicity $\frac{d(d-1)}{2}$ d1d-1 $\frac{d(d-3)}{2}$ 11 11 d1d-1
Table 1. To $O(d^{-1/2})$-order, the Hessian spectrum at type I and type II minima is identical.
Attempting to argue about distinctive analytic properties of hidden minima using the loss Hessian spectrum, one finds that eigenvalues for both types agree modulo $O(d^{-1/2})$-terms, see Table 1. In addition, in both cases the expected loss at initialization is at least an order of magnitude larger than the loss at type I and type II minima. Our investigation thus proceeds by a technique introduced in [1]. Given a critical point $C\in M(k,d)$, $d$ and $k$ fixed, we consider the functions
$m(r)$ $\coloneqq\min\{{\mathcal{L}}(W)~|~W\in\mathcal{S}_{C}(r)\},$ (1.3)
$M(r)$ $\coloneqq\max\{{\mathcal{L}}(W)~|~W\in\mathcal{S}_{C}(r)\},$ (1.4)
describing the minimum (resp. maximum) of $\mathcal{L}$} on $\mathcal{S}_{C}(r)$, the sphere of radius $r$ centered at $C$, Frobenius norm on $M(k,d)$. Of course, $m(r)$ and $M(r)$ are well-defined as $\mathcal{S}_{C}(r)$ is compact. Our approach to the pr oblem of identifying distinctive properties of hidden minima proceeds by studying various aspects of arcs $\Gamma:[0,1)\to M(k,d)$ giving $m(r)$ and $M(r)$, in particular their structure and symmetry—the focus of this work.
A formal discussion of our results requires some familiarity with the representation theory of groups and o-minimal theory. Here, we provide high-level description of our contributions, briefly covering basic definitions from group theory, and defer more detailed statements to later sections after the relevant notions have been introduced (Section 2). We consider the natural (orthogonal) action of $S_{k}\times S_{d}$ on the parameter space $M(k,d)$: the first factor permutes rows, the second columns. Given a weight matrix $W\in M(k,d)$, the largest subgroup of $S_{k}\times S_{d}$ fixing $W$, the isotropy group of $W$, is used as a means of measuring the symmetry of $W$. For example, the isotropy group of the identity matrix $I_{d}$ is the diagonal subgroup $\Delta S_{d}\coloneqq\{(\pi,\pi)~|~\pi\in S_{d}\}\subseteq S_{d}\times S_{d}$. When $T$ possesses certain invariance properties, isotropy groups occurring for minima detected empirically are seen to be symmetry breaking in the sense that they form proper subgroups of $\Delta S_{d}$. In this work, we show that arcs minimizing (1.3) or maximizing (1.4) the loss emanating from symmetry breaking critical points are themselves—symmetry breaking. More specifically,
  • We give a detailed description of the possible intersections of subspaces invariant to subgroups of $S_{d}$ (the associated isotypic components) with subspaces that are fixed by the action (Theorem 1). As tangency arcs of o-minimal definable functions must approach critical points tangentially to Hessian eigenspaces (Lemma 1), the latter amounts to obtaining an enumeration of all (generically finitely many) admissible structures and isotropy types for curves along which $C^{2}$ invariant functions are minimized and maximized. The methods apply beyond the groups considered in this work and so are of independent interest.
  • The general results are illustrated for 4 infinite families of minima of $\mathcal{L}$}: $C_{p}^{X}$, with $X\in\{\mathrm{I},\mathrm{II}\}$ denoting the family type and $p\in\{0,1\}$ isotropy $\Delta(S_{d-p}\times S_{p})$ (Theorem 2). We compute, with some effort, the two leading terms of the Hessian eigenvalues, displayed in Table 2, and show that it is precisely by these terms that the structure and symmetry of $m(r)$ associated with $\mathcal{L}$} are determined—differing, indeed, for type I and type II minima. The results cannot be obtained using existing analyses as the Hessian spectrum for type I and type II are identical to the order to which eigenvalue terms have been previously computed, see Table 1.
  • The general results are stated and proved for o-minimal structures, simplifying and generalizing existing ones for symmetry breaking. For example, the existence and construction of Puiseux series describing families of critical points follow by a direct consequence of an o-minimal version of the Curve Selection Lemma for globally subanalytic sets.
  • Curves giving $m(r)$ and $M(r)$ are instances of symmetric tangency arcs developed in [1] to show that critical points connected to symmetric ones are symmetry breaking, see concluding remarks. Here, fundamental results from o-minimal theory enable a numerical construction of tangency arcs, the continuation of which yields estimates on distances from minima, both types, to adjacent critical points. The numerical estimates conform with the theoretical analysis of the symmetry of $m(r)$ and $M(r)$.
$C^{\mathrm{I}}_{0}$ $C_{0}^{\mathrm{II}}$ $C_{1}^{\mathrm{I}}$ $C_{1}^{\mathrm{II}}$
Type I II I II
Loss $\frac{1}{\pi}+\frac{1}{2}-\frac{4}{3\pi\sqrt{d}}$ 0 $\frac{1}{\pi}+\frac{1}{2}-\frac{4}{3\pi\sqrt{d}}$ $\frac{-4+\pi^{2}}{2\pi^{2}d}-\frac{32}{3\pi^{4}d^{\frac{3}{2}}}$
Isotropy $\Delta S_{d}$ $\Delta S_{d}$ $\Delta S_{d-1}$ $\Delta S_{d-1}$
Orbit length $d!$ $d!$ $\cdot d! dd$$\cdot d$!
Rep. Mult.
$\mathfrak{t}$ 11 $\frac{\pi-2}{4\pi}+\frac{4\left(-1+\pi\right)}{\pi^{3}d}$ $\frac{\pi-2}{4\pi}+\frac{2\left(\pi-2\right)}{\pi^{2}d}$
$\frac{1}{4}+\frac{-50\pi+24+\pi^{3}+10\pi^{2}}{\pi^{4}d^{2}}$ $\frac{1}{4}+\frac{-1+2\pi}{\pi^{2}d}$
$\frac{d}{2\pi}+\frac{-\pi^{2}-4+6\pi}{4\pi\left(2-\pi\right)}$ $\frac{d}{2\pi}+\frac{-\pi^{2}-4+6\pi}{4\pi\left(2-\pi\right)}$ $\frac{d}{2\pi}+\frac{-\pi^{2}-4+6\pi}{4\pi\left(2-\pi\right)}$ $\frac{d}{2\pi}+\frac{-\pi^{2}-4+6\pi}{4\pi\left(2-\pi\right)}$
$\frac{d}{4}+\frac{1}{4}$ $\frac{d}{4}+\frac{1}{4}$
$\frac{d}{4}+\frac{-\pi^{2}-2\pi+4}{4\pi\left(2-\pi\right)}$ $\frac{d}{4}+\frac{-\pi^{2}-2\pi+4}{4\pi\left(2-\pi\right)}$ $\frac{d}{4}+\frac{-\pi^{2}-2\pi+4}{4\pi\left(2-\pi\right)}$ $\frac{d}{4}+\frac{-\pi^{2}-2\pi+4}{4\pi\left(2-\pi\right)}$
$\mathfrak{s}$ dp1d-p-1 $\frac{\pi-2}{4\pi}$ $\frac{\pi-2}{4\pi}$ $\frac{\pi-2}{4\pi}+\frac{2-\pi}{2\pi^{2}d}$ $\frac{\pi-2}{4\pi}-\frac{1}{\pi^{2}\sqrt{d}}$
$\frac{\pi-2}{4\pi}$ $\frac{\pi-2}{4\pi}+\frac{-\frac{\pi^{3}}{2}-8-\pi}{\pi^{4}d}$
$\frac{1}{4}-\frac{2}{\pi\sqrt{d}}$ $\frac{1}{4}+\frac{-1+\pi}{\pi^{2}d}$ $\frac{1}{4}+\frac{-3+2\pi}{\pi^{2}d}$ $\frac{1}{4}+\frac{-2\pi^{2}-8+7\pi}{\pi^{3}d}$
$\frac{\pi+2}{4\pi}+\frac{3\cdot\left(2-\pi\right)}{2\pi^{2}d}$ $\frac{\pi+2}{4\pi}-\frac{1}{\pi^{2}\sqrt{d}}$
$\frac{d}{4}+\frac{1}{4}$ $\frac{d}{4}+\frac{1}{4}$ $\frac{d}{4}+\frac{1}{4}$ $\frac{d}{4}+\frac{1}{4}$
$\mathfrak{x}$ $\frac{(d-p-1)(d-p-2)}{2}$ $\frac{\pi-2}{4\pi}-\frac{1}{\pi\sqrt{d}}$ $\frac{\pi-2}{4\pi}$ $\frac{\pi-2}{4\pi}-\frac{1}{\pi\sqrt{d}}$ $\frac{\pi-2}{4\pi}-\frac{1}{\pi d}$
$\mathfrak{y}$ $\frac{(d-p)(d-p-3)}{2}$ $\frac{\pi+2}{4\pi}-\frac{1}{\pi\sqrt{d}}$ $\frac{\pi+2}{4\pi}$ $\frac{\pi+2}{4\pi}-\frac{1}{\pi\sqrt{d}}$ $\frac{\pi+2}{4\pi}$
Table 2. Dominating terms of the Puiseux series describing the loss and the Hessian spectrum for 4 families of minima. The structure and symmetry of curves along which the loss is minimized is determined by the irreducible representations of $S_{d}$ with which the minimal eigenvalue (highlighted) is associated. For both types, the maximal eigenvalues (having leading terms growing linearly with $d$) belong to the $\mathfrak{t}$- and $\mathfrak{s}$-representation, implying that the dynamics concentrates in the vicinity of small subspaces of multiplicity $O(d)$.
2. Framework: the tangency set and symmetry
Focusing on the quantities $m(r)$ (1.3) and $M(r)$ (1.4) as a means of investigating hidden minima, one is naturally led to a more general consideration of curves describing critical points of ${\mathcal{L}}|_{\mathcal{S}_{\mathbf{c}}(r)}$, or equivalently curves lying in a set, the tangency set defined below, comprising all points where level sets of $f$ lie tangential to spheres centered at $\mathbf{c}$. The tangency set arises naturally in the study of singularities, for examples [37, 36, 30, 21, 38]. In this work we show that a tangency set inherits symmetries from the function with which it is associated, a result which we then use for characterizing hidden minima ([1]). A formal discussion of our main results requires some familiarity with group, representation and O-minimal theory, which we briefly review below. Proofs are deferred to the appendix.
Definition 1.
Suppose given a $C^{1}$ function $f:\mathbb{R}^{d}\to\mathbb{R}$ and a point $\mathbf{c}\in\mathbb{R}^{d}$.
  • The set of critical points of $f$ is $\Sigma(f)\coloneqq\{\mathbf{x}\in\mathbb{R}^{d}~|~Df(\mathbf{x})=0\}$.
  • The tangency set $\mho_{\mathbf{c}}(f)$ relative to $\mathbf{c}$ is defined by,
    $\mho_{c}(f)\coloneqq\{\mathbf{x}\in\mathbb{R}^{d}~|~D_{i}f(\mathbf{x})\mathbf{x}_{j}=D_{j}f(\mathbf{x})\mathbf{x}_{i},~i,j\in[d]\}.$ (2.5)
    In particular, $\Sigma(f)\subseteq\mho_{c}(f)$.
[…]
We typically do not indicate the dependence of $\Sigma$ and $\mho_{\mathbf{c}}$ on $f$ if no confusion results.
The tangency set has a particularly simple structure for quadratic forms. If $f:\mathbb{R}^{d}\to\mathbb{R}$ is given by $f(\mathbf{x})=\mathbf{x}^{\top}A\mathbf{x}/2$, $A$ symmetric, then relative to $\mathbf{c}=0$,
$\mho_{0}=\{\mathbf{x}\in\mathbb{R}^{d}~|~\exists\eta\in\mathbb{R},A\mathbf{x}=\eta\mathbf{x},~\mathbf{x}\neq 0\},$ (2.6)
the set of all eigenvectors (see Figure LABEL:fig:hc_tan for another example).
In [1], the following general result is established, relying on methods developed for the study of bifurcation phenomena in variational problems:
Theorem (Informal) If $\mathbf{c}$ is an isolated critical point, then, generically,
1. there are finitely many tangency arcs, each tangent to a Hessian eigenspace.
2. For every maximal isotropy subgroup $G\subset\Gamma$, there exists a tangency arc with isotropy $G$; if $\dim(V^{\Gamma})>0$, there exists an arc with isotropy $\Gamma$.
3. the index of the arc, defined to be the number of eigenvalues below $\langle\nabla f(\mathbf{x}),\mathbf{x}-\mathbf{c}\rangle/\|\mathbf{x}-\mathbf{c}\|^{2}$, is constant along the (open) arc. In particular, at critical points of $f$, the index of the arc and that of the Hessian coincide.
To maintain the focus of the manuscript, we do not develop the full theory presented in [1]. Instead, we restrict attention to the result above and note that a more extensive analysis is possible both globally, using for example topological methods [39], and locally, by incorporating higher-order derivatives to determine, for instance, the number of tangency arcs, their isotropy groups (not all of which need be maximal), and their indices. Below we present the results required for this paper, formulated within the framework of o-minimal structures.
O-minimal theory
The topology of tangency sets so defined can be quite complicated for arbitrary $C^{1}$ functions. However, for our applications, certain structural restrictions apply: the loss function $\mathcal{L}$}, as we show, is definable in an o-minimal structure expanding the real field [49, 50], that is, a sequence $\mathcal{D}=(\mathcal{D}_{n})_{n\in\mathbb{N}}$, $\mathcal{D}_{n}$ denoting a collection of subsets of $\mathbb{R}^{n}$, satisfying certain axioms which we detail in Section 7. Here, suffices it to note that sets defined by first-order formulae ranging over definable sets, i.e., sets belonging to $\mathcal{D}$, are themselves definable. A map $F:A\to\mathbb{R}^{n}$ is called definable if its graph $\Gamma(F)=\{(x,F(x))~|~x\in A\}$ is definable. Thus, if $f:\mathbb{R}^{d}\to\mathbb{R}$ is definable, the set $X_{\mathbf{c}}^{m}\coloneqq\cup_{r>0}\operatorname*{arg\,min}f|_{\mathcal{S}_{\mathbf{c}}(r)}$ consisting of all points at which $m(r)=\min\{f(\mathbf{x})~|~\mathbf{x}\in\mathcal{S}_{r}(\mathbf{c})\}$ (abusing notation in (1.3)) is attained, may be equivalently given by
$\{\mathbf{x}\in\mathbb{R}^{d}~|~\exists r\in\mathbb{R},\forall\mathbf{y}\in\mathbb{R}^{d}(\|\mathbf{y}-\mathbf{x}\|^{2}=r^{2}\implies\exists s\in\mathbb{R},f(\mathbf{y})-f(\mathbf{x})=s^{2})\}$ (2.7)
and so is definable (leaving the validation of use of abbreviations such as $f(\mathbf{x})-f(\mathbf{y})=s^{2}$ to the reader). Throughout, all sets and mappings involved in our analysis are definable, assuming $f$ is, as we may as $\mathcal{L}$} is (most of what we do also applies to smooth stratified mappings [34]). The proofs are straightforward and so are omitted. A prototypical example of an o-minimal structure is given by semi-algebraic sets. The loss function $\mathcal{L}$} is definable in the larger o-minimal structure $\mathbb{R}_{an}$ of globally subanalytic sets given by inverse images of subanalytic sets [33] under the Nash (algebraic and real analytic) map $\mathcal{V}_{d}(\mathbf{x})\coloneqq\left({x_{1}}/{\sqrt{1+\|\mathbf{x}\|^{2}}},\dots,{x_{d}}/{\sqrt{1+\|\mathbf{x}\|^{2}}}\right)$ mapping $\mathbb{R}^{d}$ isomorphically onto $(-1,1)^{d}$, see Section 7.
O-minimal structures, motivated as a candidate for Grothendieck’s idea of ‘tame topology’ [26], offer a framework flexible enough to carry out geometrical and topological constructions on real Euclidean set (e.g., projection) and real functions (e.g., composition and differentiation), yet sufficiently restrictive to impose certain regularity, as demonstrated by the following o-minimal theoretic result (or a metric version thereof).
Curve Selection Lemma (CSL)
If $a\in\overline{X}$, where $X$ is definable, then there exists a definable continuous map $\gamma:[0,1)\to X$ such that $\gamma((0,1))\subseteq X$ and $\|\gamma(r)-a\|=r$.
By a direct Lagrangian multipliers argument, $\mathbf{c}$ lies in the closure of $X_{\mathbf{c}}^{m}$ (definable, by (2.7)). Thus, by the CSL,
Corollary 1.
If $f$ is definable then there exists a tangency arc $\gamma$ parameterized by arc length satisfying ${\mathcal{L}}(\gamma(r))=m(r)$, and similarly for $M(r)$.
Despite the simplicity of the proof, the existence of an arc giving $m(r)$ is certainly not obvious. Counter-examples exist already for (necessarily non-definable) functions on $\mathbb{R}$, e.g., $x\mapsto\sigma^{7}(x)\sin(1/x)$. For quadratic functions, see (2.6), a tangency arc giving $m_{d}(r)$ must lie in the eigenspace associated to the minimal eigenvalue, therefore so does $\dot{\gamma}(0)$, if exists. More generally, we have the following lemma,
Lemma 1.
Suppose given a $C^{2}$ definable function $f:U\to\mathbb{R}$, $U$ open, and a critical point $\mathbf{c}\in U$. Any tangency arc parameterized by arc length approaching $\mathbf{c}$ must do so tangentially to an eigenspace of $\nabla^{2}f(\mathbf{c})$ (in particular, $\dot{\gamma}(0)$, as a one-sided limit, exists). Moreover, tangency arcs giving $m(r)$ (resp. $M(r)$) are tangential to the eigenspace associated to the minimal (resp. maximal) eigenvalue.
The lemma is proved using standard results from perturbation theory and the monotonicity theorem, an elementary result in o-minimal theory. The correspondence stated between tangency arcs and eigenspaces (see Figure LABEL:fig:hc_tan in the concluding remarks for illustration) becomes particularly useful for invariant functions on account of the detailed information on Hessian invariant subspaces holding, a-priori, for all points of fixed symmetry.
Representation theory of groups
We give a brief, if terse, review of group and representation theory that suffices for our applications. For more detail and generality see [45].
[…]
Expressing curves of critical points as Puiseux series in $d$
The loss function $\mathcal{L}$} is not only structurally tame, being $\mathbb{R}_{an}$-definable, but also has an important geometric characteristic: functionally, it only depends on inner products between the rows of $W$ and $T$, disregarding for the present weights belonging to the second layer to make the arguments below more transparent. Thus, for natural sequences (notation as in the preceding section), we may express $\mathcal{L}$} as a (definable) function of polynomials in $d$ and ${\boldsymbol{\xi}}\in\mathbb{R}^{N}$, and so regard $d$ as a real variable and so have the set $\sigma_{\mathcal{L}}\coloneqq\{({\boldsymbol{\xi}},d)\in\mathbb{R}^{N+1}~|~\Xi({\boldsymbol{\xi}})\in\Sigma({\mathcal{L}}(\cdot;d)),~d\geq d_{0}\}$ $\mathbb{R}_{an}$-definable, $d_{0}\in\mathbb{N}$ suitably chosen. Each infinite sequence of critical points $W_{d}=\Xi({\boldsymbol{\xi}}_{d}),d\in\{d_{0},d_{0}+1,\dots\}$ and $({\boldsymbol{\xi}}_{d},d)\in\sigma_{\mathcal{L}}$ gives a point in $\overline{\mathcal{V}_{N+1}(\sigma_{\mathcal{L}})}\cap\partial[-1,1]^{N+1}$ which, by the CSL, may be approached by a definable arc lying in $\mathcal{V}_{N+1}(\sigma_{\mathcal{L}})$, giving in turn a definable curve of critical points in $M(k,d)$. In $\mathbb{R}_{an}$, the entries of the resulting curve are in fact Puiseux series (in $1/d$), the coefficients of which may be computed in an obvious gradual manner, see Section LABEL:sec:puiseux.
Another important implication of $\sigma_{L}$ being definable is that by local triviality the topological type of $\Xi^{-1}\Sigma({\mathcal{L}}(\cdot;d))$ is identical for sufficiently large $d$. A continuous definable map $f:E\to B$ between definable sets is definably trivial if there are a definable set $F$, the fiber, and a definable homeomorphism $h:E\to B\times F$ such that $f\circ h^{-1}=\pi_{1},~\pi_{1}:B\times F\to B$ being the projection on the first factor. Definable maps are locally trivial in the sense that the set $B$ can partitioned into definable sets $B_{1},\dots,B_{n}$ such that $f|_{f^{-1}(B_{i})},i\in[n],$ are definably trivial. The assertion above now follows by considering the projection of $\sigma_{L}$ on the $d$-coordinate. In addition, we see that arcs of critical points may bifurcate, as they sometimes do, when $d$ is not sufficiently large.
3. Main results: structure and symmetry of tangency arcs
The isotropy groups that correspond to type I and type II minima $k=d$ (Definition 2 below) are diagonal subgroups of the form $\Delta(S_{i_{1}}\times\cdots\times S_{i_{q}})\coloneqq\{(g,g)\mbox{$\;|\;$}g\in S_{i_{1}}\times\cdots\times S_{i_{p}}\subseteq S_{d}\}\subseteq S_{d}\times S_{d},~i_{1}+\cdots+i_{q}=d$. Clearly, $S_{i_{1}}\times\cdots\times S_{i_{p}}\cong\Delta(S_{i_{1}}\times\cdots\times S_{i_{q}})$. To indicate how results given in this section are obtained, consider the orthogonal direct sum decomposition /
$M(d,d)=\mathbb{D}_{d}\oplus\mathbb{S}_{d}\oplus\mathbb{A}_{d},$ (3.8)
with $\mathbb{D}_{d}$ denoting the space of diagonal $d\times d$ matrices, $\mathbb{A}_{d}$ the space of skew-symmetric $d\times d$ matrices and $\mathbb{S}_{d}$ the space of symmetric $d\times d$ matrices with diagonal entries zero. Since $S_{d}$ acts diagonally on $M(d,d)$, the factors are $S_{d}$-invariant. We may now ask: relative to the ascending series of subgroups of $\Delta S_{d}$,
$1=\Delta S_{1}^{d}\leq\cdots\leq\Delta(S_{2}\times S_{1}^{d-2})\leq\Delta(S_{d-1}\times S_{1})\leq\Delta S_{d},$ (3.9)
what is the maximal $i\in\{0,1,\dots,d\}$ such that $M(d,d)^{\Delta(S_{i}\times S_{1}^{d-i})}$ intersects a given factor in (3.8)? Computing, we see that whereas $M(d,d)^{\Delta S_{d}}\cap\mathbb{D}_{d}=\xi I_{d}$ and $M(d,d)^{\Delta S_{d}}\cap\mathbb{S}_{d}=\xi(\mathbf{1}\mathbf{1}^{\top}-I_{d})$, the third factor $\mathbb{A}_{d}$ does not intersect $M(d,d)^{\Delta S_{d}}$. However, $\mathbb{A}_{d}$ does intersect $M(d,d)^{\Delta(S_{d-1}\times S_{1})}$. Thus, if $\mathbb{A}_{d}$ were an eigenspace of a given linear transformation then, referring to (3.9), any associated eigenvector would have isotropy type at most $\Delta(S_{d-1}\times S_{1})$. For type I and type II points, $\mathbb{A}_{d}$ is not an eigenspace. However, a similar reasoning, involving isotypic components rather than eigenspaces, applies. The isotypic decomposition of $(M(d,d),S_{d})$ is relatively simple and uses just 4 irreducible representations of $S_{d}$, each associated to a partition of the set $[d]$: the trivial representation $\mathfrak{t}$ of degree 1 associated to the partition $(d)$, the standard representation $\mathfrak{s}_{d}$ of degree d1d-1 associated to (d1,1)(d-1,1), the exterior square representation $\mathfrak{x}_{d}=\wedge^{2}\mathfrak{s}_{d}$ of degree (d1)(d2)/2{(d-1)(d-2)}/{2} associated to (d2,1,1)(d-2,1,1), and a representation $\mathfrak{y}_{d}$ of degree d(d3)/2{d(d-3)}/{2} associated to (d2,2)(d-2,2).
Theorem 1.
Write $M(d,d)=\mathbb{V}_{\mathfrak{t}}\oplus\mathbb{V}_{\mathfrak{s}}\oplus\mathbb{V}_{\mathfrak{x}}\oplus\mathbb{V}_{\mathfrak{y}}$, with the factors respectively denoting the trivial, standard, exterior square and $\mathfrak{y}$ isotypic component. If $A$ is an $S_{d}$-map, then,
  • $\mathbb{V}_{\mathfrak{t}}=M(d,d)^{\Delta S_{d}}$. In particular, referring to (3.9), the maximal isotropy type of matrices in $\mathbb{V}_{\mathfrak{t}}$ is $\Delta S_{d}$. The spectrum of $A|_{\mathbb{V}_{\mathfrak{t}}}$ and $A|_{\mathbb{V}_{\mathfrak{t}}\cap M(d,d)^{\Delta S_{d}}}$ are identical.
  • In $\mathbb{V}_{\mathfrak{s}}$, the maximal isotropy type is $\Delta(S_{d-1}\times S_{1})$. Every eigenvalue of $A|_{\mathbb{V}_{\mathfrak{s}}}$, necessarily of multiplicity $i(d-1),i\in[3]$, is an eigenvalue of $A|_{\mathbb{V}_{\mathfrak{s}}\cap M(d,d)^{\Delta(S_{d-1}\times S_{1})}}$ of multiplicity $i$.
  • In $\mathbb{V}_{\mathfrak{x}}$, the maximal isotropy type is $\Delta(S_{d-2}\times S_{1}^{2})$. The map $A|_{\mathbb{V}_{\mathfrak{x}}\cap M(d,d)^{\Delta(S_{d-2}\times S_{1}^{2})}}$ has a single eigenvalue, multiplicity one, given by the (distinct) single eigenvalue of $A|_{\mathbb{V}_{\mathfrak{x}}}$, multiplicity (d1)(d2)/2(d-1)(d-2)/2.
  • In $\mathbb{V}_{\mathfrak{y}}$, the maximal isotropy type is $\Delta(S_{d-2}\times S_{1}^{2})$. The map $A|_{\mathbb{V}_{\mathfrak{y}}\cap M(d,d)^{\Delta(S_{d-2}\times S_{1}^{2})}}$ has a single eigenvalue, multiplicity one, given by the single eigenvalue of $A|_{\mathbb{V}_{\mathfrak{y}}}$, multiplicity d(d3)/2d(d-3)/2.
A detailed description of the intersections of the subspaces involved is provided in the proof of the theorem in Section 9.
Combined with Lemma 1, we obtain the following general result quantifying the amount of symmetry breaking needed for realizing extremal tangency arcs.
Corollary 2.
(Notation and assumptions as above.) Suppose $C\in M(d,d)$ is a critical point with isotropy $\Delta S_{d}$ of an $S_{d}$-invariant definable function, and $\gamma$ a tangency arc tangential to an $\mu$-eigenspace of $\nabla^{2}f(C)$. If $\mu$ belongs to $\nabla^{2}f(C)|_{\mathbb{V}_{\mathfrak{t}}}$, then the maximal isotropy type of $\dot{\gamma}(0)$ is $\Delta S_{d}$. For $\nabla^{2}f(C)|_{\mathbb{V}_{\mathfrak{s}}}$ (resp. $\nabla^{2}f(C)|_{\mathbb{V}_{\mathfrak{x}}}$ or $\nabla^{2}f(C)|_{\mathbb{V}_{\mathfrak{y}}}$), the maximal isotropy type is $\Delta(S_{d-1}\times S_{1})$ (resp. $\Delta(S_{d-2}\times S_{1}^{2})$).
The related measure quantifying the minimal isotropy type occurring for a given isotypic component plays a major role in the study of non-local aspects of symmetric tangency sets concerning, in particular, symmetry of critical points [1], and see concluding remarks.
3.1. Structure and symmetry of tangency arcs of type I and type II minima
We now apply the general results to type I and II minima.
Definition 2.
Let $p\geq 0$ and take $G_{d}=\Delta(S_{d-p}\times S_{p})$. A family of critical points with isotropy $(G_{d})_{d\geq d_{0}}$ is type I (resp. type II) if as $d{\rightarrow}\infty$, the diagonal elements of the $(d-p)\times(d-p)$-block corresponding to the action of $\Delta S_{d-p}$ converge to 1-1 (resp. +1+1).
Below, we emphasize families with isotropy $\Delta(S_{d-p}\times S_{p}),~p\in\{0,1\}$ and $k=d$. Methods and results apply to other values of $p$, as well as for $k>d$. By Corollary 2, adapted to $\Delta(S_{d-p}\times S_{p})$-maps, we have,
Theorem 2.
The Hessian eigenvalues of type I and type II minima isotropy $\Delta(S_{d-p}\times S_{p}),p\in\{0,1\}$ represented by Puiseux series computed to two leading terms is given in Table 2. In particular, identifying the isotypic components giving the minimal and maximal eigenvalues, we find that,
  1. A.
    The maximal isotropy type of tangency arcs giving $m(r)$ for type I (resp. II) is $\Delta(S_{d-p-2}\times S_{p+2})$ (resp. $\Delta(S_{d-p-1}\times S_{p+1})$).
  2. B.
    The maximal isotropy type of tangency arcs giving $M(r)$ is at least $\Delta(S_{d-p}\times S_{p})$.
Assertions A and B, obtained by pure group representation-theoretic considerations, are easily verified by considering small values of $r>0$ using simple numerical procedures such as projected gradient descent. In the next section larger values of $r$ for $m(r)$ are considered. We note that minimal eigenvalues are also responsible for determining the isotypic component along which minima are created [7, 10].
4. Numerical results: bounding the distance to the nearest critical point
Our analysis thus far concerned local aspects of tangency arcs, specifically their symmetry and structure in the vicinity of a critical point $\mathbf{c}$. However, tangency arcs can be extended until a singularity is hit and so detect adjacent critical points. In fact, since the tangency set is definable, local triviality implies that any critical point sufficiently close to $\mathbf{c}$ may be connected to by a tangency arc. As with the metric version of the CSL, we consider the Euclidean norm $\|\cdot-\mathbf{c}\|:\mho_{\mathbf{c}}\to\mathbb{R}$. By local triviality, there exists an interval $(0,\varepsilon)$, a definable set $F$ and a definable homeomorphism $h:\mho_{\mathbf{c}}\cap\mathring{B_{\varepsilon}}(\mathbf{c})\to(0,\varepsilon)\times F$. Recalling that $\Sigma\subseteq\mho_{\mathbf{c}}$, if $\mathbf{c}^{\prime}$ is a critical point at distance at most $\varepsilon$ from $\mathbf{c}$ then $\mathbf{c}^{\prime}=h^{-1}(\varepsilon^{\prime},\mathbf{x}^{\prime})$ for some $\varepsilon^{\prime}\in(0,\varepsilon)$ and $\mathbf{x}^{\prime}\in F$. We may now define a tangency arc $\gamma(t)=h^{-1}(t\varepsilon^{\prime},\mathbf{x})$ $(\gamma(0)\coloneqq\mathbf{c})$ connecting $\mathbf{c}$ to $\mathbf{c}^{\prime}$. The set $F$ being definable has a finite number of definable (in fact, piecewise $C^{1}$-path) connected components. As a simple consequence, in each component, $t\mapsto f(\gamma(t)),t\geq 0,$ is identical for any tangency arc parameterized by arc length, whence regarding the number of tangency arcs being essentially finite (see [37] for similar results for germs of real polynomials). All of this indicates a simple effective means of estimating distances to critical points adjacent to $\mathbf{c}$: construct tangency arcs parameterized by arc length, numerically. We note that tangency arcs may also be given by analytic expressions, as with tensor decomposition problems [11].
In practice, we construct tangency arcs for $\mho_{C}({\mathcal{L}})$, $C\in M(d,d)$ a critical point, using a Lagrangian function encoding a norm constraint, $Q(W,\eta,r)={\mathcal{L}}(W)+\eta(\|W-C\|^{2}-r^{2})$. The (augmented) tangency arcs $(W(r),\eta(r))$ are computed by solving $DQ=0$ using Newton-Raphson method and small increments of $r$, typically 1e-3, until hitting a singular Jacobian or until $r_{\text{max}}$ has been reached. The former case indicates the presence of a critical point or an arc bifurcating (both occur). The latter case is taken to indicate our (very) finite approximation for an arc continuing indefinitely. The limit $r_{\text{max}}$ has been chosen so as to be an order of magnitude larger than the typical length of arcs terminating at ‘finite’ time. Following Lemma 1, arcs are initialized by setting $W(0)\coloneqq C+r_{\text{min}}B$ where $B$ denotes an eigenvector and $r_{\text{min}}=10^{-7}$. Repeating this process for different minima and ranges of eigenvalues, we found that arcs corresponding to eigenvectors associated to small eigenvalues tend to be finite and generally terminate earlier than those corresponding to large eigenvalues. In Table 4, we report the radius of arcs corresponding to minimal eigenvalues. The numerical estimates are consistent with Theorem 2:
  1. (1)
    For $C_{0}^{I}$ and $C_{1}^{II}$, the symmetry of an arc tangency realizing $m(r)$ is $\Delta(S_{d-2}\times S_{1}^{2})$ (Theorem 2.I.) and so when the ambient space $M(d,d)^{\Delta(S_{d-1}\times S_{1})}$ is replaced with the larger space $M(d,d)^{\Delta(S_{d-2}\times S_{1}^{2})}$, radii drop as expected.
  2. (2)
    For $C_{1}^{I}$, the symmetry of an arc tangency realizing $m(r)$ is $\Delta(S_{d-3}\times S_{1}^{3})$ (Theorem 2.II.) and so replacing the ambient space $M(d,d)^{\Delta(S_{d-1}\times S_{1})}$ with the larger space $M(d,d)^{\Delta(S_{d-2}\times S_{1}^{2})}$ shows no significant change. At $M(d,d)^{\Delta(S_{d-3}\times S_{1}^{3})}$, radii drop.
The results provided also suggest that for small $d$ the distance to the nearest critical point is small for spurious minima ($C_{0}^{I},C_{1}^{I}$ and $C_{1}^{II}$) compared to that of the global minima $C_{0}^{II}$. The situation changes when $d$ increases. Additional phenomena arising from the numerical study of the tangency set are a topic of current work. Non-local aspects concerning symmetry of critical points as well as tangency arcs of minimal isotropy, are studied in some depth in [1].
Ambient space $d$ $C_{0}^{I}$ $C_{0}^{II}$ $C_{1}^{I}$ $C_{1}^{II}$
$M(d,d)^{\Delta(S_{d-1}\times S_{1})}$ 7 1.16 1.16 1.25 0.90
20 1.05 1.05 1.1 0.90
100 1.01 1.01 1.03 0.97
$M(d,d)^{\Delta(S_{d-2}\times S_{1}^{2})}$ 7 0.62 1.75 1.12 0.31
20 1.01 1.53 1.08 0.81
100 1.05 1.01 1.05 0.99
$M(d,d)^{\Delta(S_{d-3}\times S_{1}^{3})}$ 7 0.62 $\infty$ 0.29 0.31
20 1.01 1.45 1.01 0.81
100 1.42 1.05 0.99
5. Concluding remarks and future work
The focus in this paper has been on $m(r)$ and $M(r)$, giving respectively the minimal and maximal value of the loss function in the vicinity of critical points. The functions are studied as a means of identifying analytic properties differentiating hidden minima, type I, from ones detected by standard gradient-based methods, type II. We prove general results on tangency arcs realizing $m(r)$ and $M(r)$ showing how pure group representation-theoretic considerations yield a precise description of the admissible types of such arcs. The general results used for the loss function $\mathcal{L}$} reveal that tangency arcs of type I differ from type II by their structure and symmetry, provably requiring a greater extent of symmetry breaking to have curves realizing $m(r)$ on account of $O(d^{-1/2})$-terms of the Hessian spectrum of type I and II which are otherwise identical. The theoretical results are derived using methods developed for o-minimal structures which imply in particular a certain topological regularity required for the numerical work presented. In addition to confirming the structure and symmetry type predicted by our theoretical results, the construction of tangency arcs provides an effective means for studying critical points adjacent to a given one.
Bias terms
A natural question arising from the analysis of minima in this work is how the conclusions drawn in the unbiased setting are affected by the introduction of bias terms. The biased analogue of the kernel appearing in the unbiased analysis is as follows. For $\mathbf{w},\mathbf{v}\in\mathbb{R}^{d}$ and $a,b\in\mathbb{R}$, define
$k(\mathbf{w},a,\mathbf{v},b):=\mathbb{E}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},I_{d})}\!\left[\sigma(\langle\mathbf{w},\mathbf{x}\rangle+a)\,\sigma(\langle\mathbf{v},\mathbf{x}\rangle+b)\right].$
Assume $\|w\|>0$ and $\|v\|>0$ (the edge cases follow similarly). Set $\rho=\langle\mathbf{w},\mathbf{v}\rangle/(\|\mathbf{w}\|\|\mathbf{v}\|)$, $\alpha=-a/\|\mathbf{w}\|$, $\beta=-b/\|\mathbf{v}\|$, and $s=\sqrt{1-\rho^{2}}$. Let $\phi$ and $\Phi$ denote the standard normal density and distribution functions, and let $\Phi_{2}(\cdot,\cdot;\rho)$ denote the bivariate normal distribution function with correlation $\rho$. Define also the bivariate normal density
$\phi_{2}(u,v;\rho)=\frac{1}{2\pi\sqrt{1-\rho^{2}}}\exp\!\left(-\frac{u^{2}-2\rho uv+v^{2}}{2(1-\rho^{2})}\right).$
Then the kernel admits the explicit expression
$k(\mathbf{w},a,\mathbf{v},b)={}$ $\ \|\mathbf{w}\|\,\|\mathbf{v}\|\Bigg(\rho\Big[1-\Phi(\alpha)-\Phi(\beta)+\Phi_{2}(\alpha,\beta;\rho)\Big]+(1-\rho^{2})\,\phi_{2}(\alpha,\beta;\rho)\Bigg)$
$\quad+\ \|\mathbf{w}\|\,b\;\phi(\alpha)\Big(1-\Phi\!\big(\tfrac{\beta-\rho\alpha}{s}\big)\Big)+\ \|\mathbf{v}\|\,a\;\phi(\beta)\Big(1-\Phi\!\big(\tfrac{\alpha-\rho\beta}{s}\big)\Big)$
$\quad+\ ab\Big[1-\Phi(\alpha)-\Phi(\beta)+\Phi_{2}(\alpha,\beta;\rho)\Big].$
Remark 2 (Owen’s $T$-function).
The bivariate normal distribution function $\Phi_{2}$ can also be expressed using Owen’s $T$-function, which yields equivalent one-dimensional integral representations that are often employed for numerical evaluation.
First- and second-order derivatives of the kernel likewise admit explicit expressions. For example, the gradient with respect to $\mathbf{w}$ takes the form
$\nabla_{\mathbf{w}}k(\mathbf{w},a,\mathbf{v},b)=c_{1}\,\mathbf{w}+c_{2}\,\mathbf{v},$
where
$\Delta$ $\coloneqq\langle\mathbf{w},\mathbf{w}\rangle\langle\mathbf{v},\mathbf{v}\rangle-\langle\mathbf{w},\mathbf{v}\rangle^{2}=\|\mathbf{w}\|^{2}\|\mathbf{v}\|^{2}(1-\rho^{2}),$
$c_{1}$ $=\frac{\langle\mathbf{v},\mathbf{v}\rangle\,\Gamma_{1}-\langle\mathbf{w},\mathbf{v}\rangle\,\Gamma_{2}}{\Delta},\qquad c_{2}=\frac{-\langle\mathbf{w},\mathbf{v}\rangle\,\Gamma_{1}+\langle\mathbf{w},\mathbf{w}\rangle\,\Gamma_{2}}{\Delta}.$
$\Gamma_{1}$ $=\|\mathbf{w}\|\Bigg(\|\mathbf{v}\|\Big(\rho\big[1-\Phi(\alpha)-\Phi(\beta)+\Phi_{2}(\alpha,\beta;\rho)\big]+(1-\rho^{2})\phi_{2}(\alpha,\beta;\rho)\Big)$
$+b\,\phi(\alpha)\Big(1-\Phi\!\big(\tfrac{\beta-\rho\alpha}{s}\big)\Big)\Bigg),$
$\Gamma_{2}$ $=\|\mathbf{v}\|\Bigg(\|\mathbf{v}\|\Big(1-\Phi(\alpha)-\Phi(\beta)+\Phi_{2}(\alpha,\beta;\rho)+(1-\rho^{2})\,\beta\,\phi(\beta)\Big(1-\Phi\!\big(\tfrac{\alpha-\rho\beta}{s}\big)\Big)$
$+\rho(1-\rho^{2})\phi_{2}(\alpha,\beta;\rho)\Big)+b\,\phi(\beta)\Big(1-\Phi\!\big(\tfrac{\alpha-\rho\beta}{s}\big)\Big)\Bigg).$
A detailed and systematic analysis of the structural and symmetry properties, together with precise asymptotic characterizations of the eigenspectrum, extending the methodology developed in the present work, and a comparative study of the resulting classes of global and spurious minima associated with the biased kernel, is carried out in [2].
Symmetry breaking and deep architectures
In general terms, this article developed out of a program to understand how symmetry breaking phenomena exhibited by nonconvex loss landscapes coming from natural distributions may allow gradient-based methods to find good minima efficiently. The ability to control the extent to which tangency arcs break the symmetry of critical points as demonstrated in the paper is an instance of a general approach developed in [1] for arguing about tractability of symmetric nonconvex optimization problems.
Critical points of $G$-invariant functions may or may not be symmetric. In [1], it is shown, however, that when certain conditions apply, critical points connected to symmetric ones by tangency arcs are symmetry breaking. Once a lower bound on the isotropy of critical points has been obtained, the effects of increasing the number of neurons on spurious minima and, crucially, on the possible emergence of descent directions can be analyzed, e.g., [9].
A rigorous and systematic analysis of these phenomena in loss landscapes arising from the training of deep neural architectures is developed in [3]. In that work, the theory of distributions is adopted as a comprehensive analytical framework that permits a precise treatment of the nonsmoothness of ReLU networks. Within this framework, tools from geometric measure theory are employed to obtain a refined characterization of the gradient, interpreted as a vector-valued function of bounded variation, and of the Hessian, which is accordingly represented by Radon measures.
6. Acknowledgements and disclosure of funding
We thank Tierra del Sol. The research was supported by the Israel Science Foundation (grant No. 724/22).